A bond with 14 years to maturity and a coupon rate of 7​% has a​ par, or​ face, value of $22,000. Interest is paid annually. If you require a return of 11​% on this​ bond, what is the price of the​ bond? ​(Round to the nearest​ cent.)
Answer #1

Using formula for annual payments,

BP= C * (( 1 - ( 1 + r )^(-t) ) / r ) + F / ( 1 + r )^t

where, BP = Bond Price

r = rate of interest

C = Coupon Price

t = time in years

here,

r = 11% = 0.11

time = t = 14

F = face value = $22000

C = 7% of Face value = 7 * 220 = $1540

BP= 1540 * (( 1 - ( 1 + 0.11 )^(-14) ) / 0.11 ) + 22000 / ( 1 + 0.11 )^14 = 10752.072453 + 5103.886 = $15855.958

Hence, the bond price is $15855.958